OpenStudy (anonymous):

A wooden pyramid, 12 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 5 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids.

OpenStudy (anonymous):

@agentx5

OpenStudy (anonymous):

 \[ \frac{b}{h^2} \int_0^h (h-y)^2 \, dy = \frac{-b}{3h^2} (h-y)^3 \bigg|_0^h = \tfrac{1}{3}bh. \] V = (1/3)(base)(height) Ratio = firstVolume:secondVolume